Find the no. of atom in NH3

Grade:11

3 Answers

Aarti Gupta
8 years ago
NH3 or ammonia molecule is formed of nitrogen and hydrogen atoms.Its an covalent compound formed by shairing of electrons between one nitogen atom and three hydrogen atoms.So we can say that 1molecule of ammonia consist of 1N atom and 3H atoms,so there are 4atoms in this compound.
rutvik
8 years ago
it s easy anuj manhas
 
 molar mass of N = 14
molar mass of h = 3 becoz.. there 3 hydrogen molecules or atoms
therefore NH3 = 14 * 3
                     = 42
Now,
1 g-atom NH= 42g
therefore g-atoms * N= no. of atoms
42 * 6.022 * 1023  = no. of atoms
252.92 * 1023  = no. of atoms
 
Hence, this is solution so simple 
Rahul Jiji George
8 years ago
1 moleculeof amminoa consists of 4 atoms
i mole of ammonia consists of 4 * 6.022*10^23 atoms